OpenStudy (anonymous):

What is the sum of the geometric sequence 1, -6, 36, … if there are 7 terms?

OpenStudy (amistre64):

6S = 6, -36, 216, ... r^(n-1), -r^n +S = 1, -6, 36, -216, ...-r^(n-1) ----------------------------- 7S = 1 - r^n S = (1-r^n)/7 r = -6, n=7

OpenStudy (anonymous):

so it would be 6665

OpenStudy (amistre64):

in general; 1-r^n ----- ; is the sum of a geometric series 1-r

OpenStudy (anonymous):

so if it was out of 6 terms instead would my answer just be -6665 or 39991

OpenStudy (amistre64):

hmm, id have to brute it to be sure :) the r=-6 is throwing me off

OpenStudy (anonymous):

theses are the 4 questions---What is the sum of the geometric sequence -1, 6, -36, … if there are 7 terms? A) -39,991 B) -6,665 C) 6,665 D) 39,991 26. What is the sum of the geometric sequence -1, 6, -36, … if there are 6 terms? A) -39,991 B) 6,665 C) -6,665 D) 39,991 27. What is the sum of the geometric sequence 1, -6, 36, … if there are 6 terms? A) -39,991 B) 6,665 C) -6,665 D) 39,991 28. What is the sum of the geometric sequence 1, -6, 36, … if there are 7 terms? A) -39,991 B) -6,665 C) 6,665 D) 39,991

OpenStudy (amistre64):

1 S = (-6)^0 +(-6)^1 +(-6)^2 +(-6)^3 +(-6)^4 +(-6)^5 -(-6)S = -(-6)^1 -(-6)^2 -(-6)^3 -(-6)^4 -(-6)^5 - (-6)^6 ----------------------------------------------------------- 7S = 1 - (-6)^6 S = (1-(-6)^6)/7

OpenStudy (amistre64):

so the geometric seq is the same, all that changes are the number of terms

OpenStudy (amistre64):

n is the number of terms; yes

OpenStudy (anonymous):

so wld the answers be a,b,d,b

OpenStudy (amistre64):

if thats what the calculator pops out, then yes

OpenStudy (amistre64):

the 3rd one is the negative of the 2nd one; so thats C i believe

OpenStudy (amistre64):

and the last is the negative of the first; so D on that

OpenStudy (anonymous):

thanks soo muchhh

OpenStudy (amistre64):

yw